# Accurate Unsupervised Photon Counting from Transition Edge Sensor Signals

### v2

@@ VI Conclusion
Nonlinear methods like t-SNE and UMAP that aim to preserve local data structures offer better resolution over photon numbers in the case of transition edge sensor signals compared to previously used techniques like signal area and PCA. These methods can be used directly to replace currently used methods, with the caveat that they cannot predict new samples without computing the entire dataset.
[…]
Beyond TES devices, the techniques explored in this work hold promise for enhancing the performance of other single-photon detectors, such as SNSPDs. For instance, principal component analysis (PCA) has shown potential in processing SNSPD signals [46, 24], highlighting the versatility of these approaches across photon-detection technologies.
@@ Acknowledgements
Appendix A Neural network
For both parametric implementations of t-SNE and UMAP, we use a simple feed-forward neural network defined as a series of blocks containing linear layers with ReLU activation functions followed by a batch normalization step. The results presented in this work use a neural network with 4 blocks, where each linear layer contains 300 inputs and outputs. While more complex architectures could be used for this task, we find that even an elementary neural network can achieve this task accurately, resulting in fast data transformation.
We note that we use the same neural network to predict data for both close-to-uniform and close-to-geometric cases. This is done to train the neural network on a balanced dataset, and in this process we guarantee that the neural network is never trained on test data. Using different distributions for the training step is not advantageous to parametric methods, since the close-to-uniform dataset contains fewer samples than in the close-to-geometric case. Additionally, other methods do not benefit from being trained using this data.

### v3

@@ VI Conclusion
Nonlinear methods like t-SNE and UMAP that aim to preserve local data structures offer better resolution over photon numbers in the case of transition-edge sensor signals compared to previously used techniques like signal area and PCA. These methods can be used directly to replace currently used methods, with the caveat that they cannot predict new samples without computing the entire dataset. However, with a large dataset (u=550 000u=550\,000 samples), we demonstrate the potential of neural networks that recreate the embedding of t-SNE and UMAP. These models remain simple and could be further explored, offering a promising direction for future research. Enhancing the generalization capabilities of these models could enable their application in real-time photon-number resolution.
This enhanced photon-number resolution has direct implications for the preparation of non-classical states of light, such as cat states, GKP states, and magic states [65], which require precise heralding and photon-number discrimination. In particular, higher-resolution detection allows for more accurate post-selection, improved state fidelity, and greater control over resource-state generation, which are all critical for fault-tolerant quantum computation and quantum error correction.
Beyond TES devices, the techniques explored in this work hold promise for enhancing the performance of other single-photon detectors, such as SNSPDs. For instance, principal component analysis (PCA) has shown potential in processing SNSPD signals [49, 26, 66], highlighting the versatility of these approaches across photon-detection technologies.
During the publication process, we became aware of an independently conducted work that similarly explores the use of advanced signal classification techniques to improve TES detection performance [67]. We acknowledge this complementary work, which highlights the growing interest in applying machine learning to enhance the speed and efficiency of photon-number resolving detectors.